An article on estimating 28-day strength of concrete considered regressing y = 28-day standard-cured strength (psi) against x = accelerated strength (psi). Suppose the equation of the true regression line is y = 1830 + 1.3x.

(a) What is the expected value of 28-day strength when accelerated strength = 2490?
psi

(b) By how much can we expect 28-day strength to change when accelerated strength increases by 1 psi?
psi

(c) By how much can we expect 28-day strength to change when accelerated strength increases by 130 psi?
psi

(d) By how much can we expect 28-day strength to change when accelerated strength decreases by 130 psi?
psi
Answer #1

regression line is
y = 1830 + 1.3x

a)
for x = 2490
y = 1830 + 1.3*2490
y = 5067

b)
strenght to change by 1.3 psi

c)
if accelerated strength is increased by 130 psi
the strength increases by 1.3 * 130 = 169 psi

d)
if accelerated strength is decreased by 130 psi
the strength decreases by 1.3 * 130 = 169 psi